OpenStudy (marinadanyel):

What is the oblique asymptote of the function f(x) = the quantity of x squared plus 7 x plus 11, all over x plus 5 ?

OpenStudy (marinadanyel):

\[\frac{ x^2+7x+11 }{ x+5 }\]

OpenStudy (marinadanyel):

@Photon336

OpenStudy (marinadanyel):

y = x + 1 y = x − 2 y = x + 4 y = x + 2

OpenStudy (photon336):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):